Please help me how to solve this problem,A water tank is to contain 6 cubic meters. If its length is 4m and its width is 2m, find the height of the tank.
Answer by sachi(548)

Please help me how to solve this problem,A water tank is to contain 6 cubic meters. If its length is 4m and its width is 2m, find the height of the tank.

ans:
length=4m width=2m
volume=length*width*height=6 cum
so height=6/(4*2)=6/8=3/4m=75 cm ans